OpenStudy (xoxolove456):

Two cars left an intersection at the same time, one heading due north and the other due east. Later the cars were exactly 125 miles apart. The car headed north had traveled 32 miles further than the car heading east. How far did each car travel???

OpenStudy (mww):

 Hope this helps. Let x be distance travelled east by B then A travelled x + 32. Use pythagoras of course and simplify to get x.

OpenStudy (xoxolove456):

Im not familiar with that

OpenStudy (mww):

Pythagoras' Theorem states the square of the hypotenuse is the sum of the other two squares \[c^2 = a^2 +b\] Here we have c being 125 (the hypotenuse), then a and b are x and x + 32. You must draw a diagram like a showed above to illustrate, otherwise you will not see this relationship! So plugging these into Pythagora's Equation \[125^2 = x^2 + (x+32)^2 = x^2 + x^2 + 64x + 32^2 = 2x^2 + 64x + 32^2\] This solves as \[2x^2 + 64x + 32^2 - 125^2 = 0 \rightarrow 2x^2 + 64x -14601 = 0\] solve that quadratic equation to 0 using Quadratic formula and then pick an x such that x > 0.